OpenStudy (anonymous):

Next question of the day by big sexy himself... Factor out polynomial completely 2x2-12x-80 = ?

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

Delaine dont reply on my pellet before i come kick you in the damn ear!

OpenStudy (anonymous):

i got (x - 20) (x + 8)

OpenStudy (anonymous):

deff wrong answer tardo nelson!

OpenStudy (anonymous):

A correct answer would be swell...

OpenStudy (anonymous):

\[2x^2-12x-80=2(x^2-6x-40)=2(x-10)(x+4)\]

OpenStudy (anonymous):

trick was to factor out the "2" first

OpenStudy (anonymous):

Thats still wrong bucko

OpenStudy (anonymous):

2 (x +4) (x -10) that should be right..

OpenStudy (anonymous):

? is there a difference between \[2(x+4)(x-10)\] and \[2(x-10)(x+4)\]?

OpenStudy (anonymous):

I didnt type the 2 i was wrong.

OpenStudy (anonymous):

*hugs*